OpenStudy (anonymous):

WILL GIVE MEDAL TO BEST RESPONSE Use substitution to determine which of the following points is a solution to the standard form equation below. 3x-5y=8 a. (1,1) b. (1,-1) c. (-1,1) d. (-1,-1)

OpenStudy (anonymous):

multiply so you can subtract to eliminate a variable (3x - 5y = 8) x4 12x - 20y = 32 (4x - 7y = 10) x3 12x - 21y = 30 subtract the second from the first... 12x - 20y = 32 - (12x - 21y = 30) y = 2 substitue the y back into the original problems 3x - 5 (2) = 8 3x - 10 = 8 3x =18 x = 6 y = 2 and x = 6 if put them both back into the problem then you should get the answer on the other side....

OpenStudy (johnweldon1993):

so it is just this equation? 3x-5y=8 Well just take each of the potential points ...and substitute them in for 'x' and for 'y' so for example....choice A says (1 , 1) (x , y) so replace ever 'x' with 1 and 'y' with 1 3x - 5y = 8 3(1) - 5(1) = 8 3 - 5 = 8 -2 = 8 This is not true...so it is not the right choice....just follow that for the other choices until you find the answer